Asher had 2635g of lithium metal. He also had a tank of nitrogen gas with a volume of 245 L, a pressure of 6000 tore, and at a temperature of 8 degrees C. How much lithium nitride could Asher make in a synthesis reaction? What is the limiting reagent?
1
Expert's answer
2019-04-11T07:47:12-0400

2Li + 3N2 = 2Li3N

n(Li) = m / M = 2635 / 7 = 376.4 (mol)

pV = nRT

n = pV/RT = (799934.2 * 0.245) / (8.314 * 281) = 83.9 (mol)

N2 is the limiting reagent

n(Li3N) = n(N2) * 2 / 3 = 55.9 (mol)

m(Li3N) = n * M = 55.9 * 34.8 = 1945.32 (g)

Answer: m(Li3N) = = 1945.32 (g), N2 is the limiting reagent